OpenStudy (anonymous):

Give the equivalent numerator if the denominator is (a + 1)³.

OpenStudy (anonymous):

\[\frac{ 3a }{ (a+1)^2 }\]

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

Well, it's the same as the previous one. We have to multiply both numerator and denominator by a expression such that it gives us the denominator of a equivalent function whose denominator is (a+1)^3

OpenStudy (owlcoffee):

That is if we have: \[\frac{ 3a }{ (a+1)^2 }\] We'll just multiply both the numerator and denominator by (a+1). It works because it is the same expression and we just miltuply it, because we can look at the (a+1)^2 like: \[\frac{ 3a }{ (a+1)(a+1) }\] So, we will multiply both numerator and denominator by (a+1): \[\frac{ 3a(a+1) }{ (a+1)(a+1)(a+1) }\] so, simplifying: \[\frac{ 3a(a+1) }{ (a+1)^3 }\] so, we do the distributive on the numerator: \[\frac{ 3a^2+3a }{ (a+1)^3 }\] And there you have it.

OpenStudy (anonymous):

3a(a + 1) 3a(a + 1)² 3a(a + 1)³

OpenStudy (anonymous):

So which one is it?

OpenStudy (owlcoffee):

take a look at the second to last expression I wrote in the last response, and the numerator should give you the answer.

OpenStudy (anonymous):

#2?

OpenStudy (owlcoffee):

No, try again